A nucleus X initially at rest , undergoes alpha decay acc. to the equation
232(X)Z--->A(Y)90 +ALPHA.[232 is mass no. ,Z is atomic no.same for Y].
What fraction of the total energy released in the decay will be the kinetic energy of the alpha particle ?
{ans:-228/232}

Grade:12

1 Answers

vikas askiitian expert
13 years ago

let E amount of energy is released in this process then this energy is converted to kinetic energy of alfa particle & daughter nuclie...

initially the system is in rest so momentam = 0

 

finally momentam = PA + PD

PA = momentam of alfa particle & PD is momentam of daughter nuclie ....

here we can apply conservation of momentam so

PA + PD = 0 ................1

total kinetic energy = (KE)A + (KE)D = E .............2

kinetic energy(KE) = p2/2m so eq 1

(KE)AMA = (KE)DMD  .....................3

from eq 3 & 2

(KE)A + (KE)AMA/MD = E

 (KE)A(MA+MD)/MD = E

(KE)A/E = MD/(MA+MD)=228/232      

this is the required fraction